# HuggingGraph: Understanding the Supply Chain of LLM Ecosystem

### v2

@@ 4. Experiments and Finding
  • RQ #4: What are the supply chain between datasets?
  • RQ #5: What are the relationships between models and datasets?
  • RQ #6: What insights can be gained from the dynamic updates?
  • RQ #7: How can HuggingGraph be applied to other platforms?
@@ 4.1. RQ #1: Supply Chain Graph Properties
Figure 3. (a) Indegree distribution and (b) Outdegree distribution. X-axis represents degrees, and the Y-axis represents the number of vertices in a logarithmic scale.
@@
Graph basics. The collected supply chain graph is a medium-scale directed heterogeneous graph with 402,654 nodes and 462,524 edges as of June 30th, 2025. In particular, there are six different types of nodes, including 28,384 base models, 115,211 fine-tuned, 79,254 adapters, 98,143 quantization models, 13,028 merges, and 68,634 datasets. The average degree is about 1.15, denoting that it is a very sparse graph. Furthermore, we identified substantial metadata missing. As of June 30, 2025, among 1.8 million models, only 50,156 (2.76%) include a model tree, while $\sim$ 550K models lack any metadata beyond their name, with nearly 400K entries empty or invalid. Similarly, of the 450K datasets, only 68,634 (15.26%) provide datacards, leaving $\sim$ 380K datasets without any usable metadata.
@@
Degree distribution of a graph describes how node degrees (the number of edges connected to a node) are distributed across the graph. Figure 3 illustrates the indegree and outdegree distribution of the graph. We show not only the total distribution but also the distribution of six types of nodes, including base models, fine-tuned models, adapter models, quantization models, merged models, and datasets. We made two interesting observations. (i) This degree distribution in our supply chain graph shows a heavy-tailed behavior. In particular, the indegree distribution shows a large spread across different categories. The outdegree distribution follows a similar pattern but may differ in specific cases (e.g., adapters seem to have a more restricted degree distribution). The heavy-tailed behavior suggests that most nodes have low degrees, while a few central nodes (hubs) dominate the graph. In particular, the dataset macrocosm-os/images has the highest indegree value of 550, and Mistral-7B-v0.1 from “mistral AI” has the highest outdegree value of 1,093. Specifically, the base models act as high-degree hub nodes as they are heavily used by other task-specific models.
@@ 4.2. RQ #2: Supply Chain Structural Analysis
Figure 4. Cumulative distribution function (CDF) of WCC size.
@@
Connectivity analysis analyzes the structural connectivity information. We performed weakly connected components (WCC), which identifies maximal sets of nodes that remain connected when edge directions are ignored. The total number of WCCs in our supply chain graph is 44,908. Figure 4 shows the cumulative distribution function (CDF) of the WCC distribution. We made two interesting observations. (i) The largest WCC covers 247,244 nodes, accounting for 61.8% of all the nodes. It reflects the dense interconnections that pervade the ecosystem. This vital element is essential for effective information sharing, resource allocation, and structural support and is the base of the ecosystem. In the largest WCC, major models are included, such as Gemma-2B, DistilBERT, and GPT-2. (ii) In contrast, the remaining WCCs collectively hold 38.6% of the nodes (i,e,.155,410 nodes), with most having 1, 2 or 3 nodes. This indicates a fragmented outer edge characterized by specialized models, rare datasets, or active experimental projects. The prevalence of these small, isolated pieces suggests niche attempts that lack integration with the overall system. For example, zhongqy/RMCBench (benchmarking dataset) and yigit69/bert-base-uncased-finetuned-rte-run_3 (recognizing textual entailment task) remain disconnected due to limited reuse or insufficient metadata.
@@
We also computed strongly connected component (SCC), a maximal subgraph in which every node is reachable from every other via directed edges, identifying 398,198 SCCs. Remarkably, only 591 of these are non‑trivial (size ¿ 1), collectively encompassing 2,169 nodes (0.54% of the graph). The largest SCC comprises 478 dataset nodes, among them tree‑of‑knowledge and OpenHermes‑2.5, forming a tightly‑knit cluster. By contrast, the remaining 99.46% of nodes each reside in trivial (size‑1) SCCs.
Community detection is the process of identifying groups or clusters of nodes within a graph that are more densely connected internally than with the rest of the graph. In the context of the LLM supply chain, it reveals semantically or functionally aligned subgraphs that reflect the patterns of reuse and task-specific models or datasets. We use the Louvain method (De Meo et al., 2011), which is a greedy optimization method for maximizing modularity. Modularity is a measure of how well a graph is divided into communities, where a high modularity score indicates that most edges fall within communities rather than between them.
@@
Table 2 summarizes the top-10 communities detected using the Louvain method. Our analysis reveals an exceptionally modular structure, with each of the top communities achieving a high modularity score of 0.96, indicative of strong intra-community connectivity. Collectively, these communities span a wide range of functional domains, underscoring the presence of well-defined, task-aligned clusters within the ecosystem. We made two interesting observations. (i) The largest community consists of 9,390 nodes and attains a modularity score of 0.96, indicating an extremely cohesive internal structure. It revolves around base models like OLMoE and CausalLM, and general-purpose datasets like prompt-perfect. This suggests a densely connected cluster facilitating widespread reuse and fine-tuning. (ii) Several other communities reflect clear task-based segmentation. For instance, Community 3 (7,388 nodes, modularity 0.96) focuses on benchmarking, with models like Wanxiang and datasets like marco-o1. Similarly, ID 5 (4,554 nodes, modularity 0.96) centers on instruction-tuning, connecting models like MedLlama-3-8B with curated datasets like dpo-mix.
@@
Finding #2: The LLM supply chain graph features a dominant core (61.8% of nodes), while high modularity (0.96) reveals task-aligned, semantically coherent communities amid a fragmented periphery.
4.3. RQ #3: Supply Chain Analysis of LLM Models
@@
Base model impact. We would like to understand the impact of base models. Here, we quantify the impact of a base model as the number of task-specific models that depend on it. The more dependencies, the larger the impact it has. We start with a base model and perform forward analysis by computing a breadth-first search (BFS) following the outgoing edges to get that from our LLM supply chain graph. This leads to a forward subgraph, which denotes all the models that depend on the base model, including fine-tuned, adapted, quantized, or merged models.
Table 3 shows the top-10 base models sorted by the forward subgraph size, which is the number of impacted task-specific models. We make two interesting observations. (i) A base model can significantly impact the LLM supply chain ecosystem. For example, Llama-3.1-8B is a base model from Meta used for efficient text generation, code assistance, and research (He et al., 2024). Due to its relatively small size, which allows for deployment in resource-constrained environments, making advanced AI accessible to broader stakeholders (Grattafiori et al., 2024). It has generated up to 7,544 models, including 1,710 fine-tuned versions, 1,542 adapters, 3,473 quantizations, and 1,693 merged models tailored to specific tasks. (ii) In terms of fine-tuning, the base model Mistral-7B-v0.1 has been fine-tuned the most times, totaling 2,105. It is a faster and lighter version of the Mistral model from Mistral AI. It is obtained by training on a large corpus with grouped-query and sliding window attention by Mistral AI, delivering efficient text generation, NLP, and code assistance on consumer hardware for low-latency applications like chatbots and text classification (Thakkar and Manimaran, 2023).
@@
Llama-3.1-8B 7544 1710 1542 3473 1693 25
Mistral-7B-v0.1 6744 2105 2187 1435 1254 27
Qwen2.5-7B 6733 1972 1764 2516 1132 11
Meta-Llama-3-8B 5633 967 1511 2220 1967 21
Llama-3.1-70B 4063 698 281 2075 2519 11
Qwen2.5-32B 3909 1086 158 2311 1049 12
Qwen2.5-1.5B 3645 1300 1290 949 248 8
Qwen2.5-0.5B 3521 1669 1006 810 46 11
Qwen2.5-14B 3362 726 411 1880 1166 15
Meta-Llama-3-8B-Instruct 3118 640 405 1394 1305 34
Task-specific model analysis. Given a task-specific model, we want to understand how it evolves, meaning what other models it has relied on. To achieve that, for each task-specific model, we perform a backward analysis by running a BFS following the incoming edges. To that end, the derived subgraph shows the models it relies on.
Table 4 shows the top 10 task-specific models sorted by the backward subgraph size, which is the number of models they rely on. We make two interesting observations. (i) A fine-tuned model, command-r-1-layer, illustrates the depth and complexity of transformations in the LLM supply chain. This model operates in bfloat16 (BF16) precision for efficient text generation and natural language understanding (Pignatelli et al., 2024), originates from the base model cc4ai, and has undergone extensive lineage evolution before reaching its final form. Specifically, it depends on 40 upstream artifacts, including 39 other fine-tuned models, and spans 39 transformation levels in its backward lineage chain, as detailed in Table 4. (ii) We observed that adapters are mainly used for lightweight fine-tuning and merges for model integration, but task-specific models like command-r-1-layer, as optimized standalone derivatives, do not evolve from adapters or merges in their backward lineage (Pignatelli et al., 2024).
4.4. RQ #4: Dataset Supply Chain Analysis
@@
This research question aims to provide comprehensive insights into the different usages of the datasets within the LLM supply chain graph. There are two relationships between datasets. On one hand, one dataset can be created by combining a variety of other datasets from disparate sources. On the other hand, one dataset can serve as a building block for new datasets.
In the LLM supply chain graph, 68,634 datasets act as training datasets that are used for model training or generating new datasets. Table 5 presents the top 10 datasets ranked by the highest number of included and derived datasets. (i) We observe that a training dataset can include many small datasets. Sitting at the top is macrocosm-os/images (Kido, 1961) from macrocosm. It is included in 550 datasets and is a general-purpose image modification corpus used in advanced multimodal AI research and large-scale vision-language model training. Moreover, bespokelabs/Bespoke (Elliot et al., 2024) from bespokelabs is part of 215 datasets, is a high-quality synthetic dataset designed to enhance multimodal AI research.
Table 5. Top-10 datasets by # of included & derived datasets.
Dataset # of included Dataset # of derived
macrocosm-os/images 550 HuggingFaceH4 989
bespokelabs/Bespoke 215 CodeFeedback 857
databricks/databricks 139 MADLAD 850
Open-Orca/OpenOrca 121 Capybara 823
nguha/legalbench 117 Glot500 804
OpenAssistant/oasst1 112 dolphin-coder 756
LDJnr/Capybara 108 SlimOrca 726
kvn420/Tenro_V4.1 105 orca_dpo_pairs 687
teknium/OpenHermes 104 gutenberg 675
Anthropic/hh-rlhf 98 samantha 658
(ii) The other observation is that a single dataset can be derived to many datasets, showing the huge overlaps between the datasets and also the models trained with them. The right two columns of Table 5 show the top 10 datasets sorted by the maximum number of derived datasets they have been included in. The leading contributor, HuggingFaceH4 from HuggingFace has been derived to 989 datasets. It is primarily used for natural language understanding and generation tasks, including chatbots, text generation, code completion, reasoning, and multilingual processing, enabling the training and enhancement of models specialized in advanced conversational and instruction-following capabilities (Dietrich and Hollstein, 2025). Similarly, the CodeFeedback dataset from m-a-p has facilitated the creation of 857 datasets, which is a collection of high-quality code instruction queries designed to enhance the training of large language models (LLMs) for code generation, debugging, and explanation tasks, enabling improved performance in complex programming scenarios(Meyers et al., 2018).
Finding #4: Datasets play critical roles in training with 68,634 datasets contributing to model development through inclusion and derivation, as seen in macrocosm-os/images and HuggingFaceH4.
4.5. RQ #5: Supply Chain between Models and Datasets
@@
This research question aims to explore the interconnections between models and datasets within the supply chain graph. It provides valuable insights from dual perspectives, including one dataset versus multiple models and one model versus multiple datasets.
@@
One dataset versus multiple models refers to the case when a single dataset is used to train multiple models. Table 6 shows the top 10 datasets based on the number of models trained on them. In particular, Mistral-7B-v0.1 takes the leading position and is a widely adopted open-source dataset known for its strong performance in general-purpose language understanding and generation tasks. It has been used to train 1,093 models, including 300 fine-tuned variants, 300 adapters, 193 quantized models, and 300 merged models, highlighting its broad adoption across diverse model derivation strategies.
In addition, the dataset TinyLlama-1.1B-v1.0, a compact and efficient model variant designed for low-resource deployment, is used to train 728 models, featuring 300 fine-tuned variants and 300 adapters. Similarly, open_llama_3b, an open-access dataset of LLaMA, supports 285 adapter-based models, indicating a preference for lightweight, modular adaptation. The dataset Yarn-Mistral-7b-128k also shows significant reuse, contributing to 279 adapters and 14 quantized models. Furthermore, MAmmoTH2-7B stands out with 210 merged models, showcasing its role in ensemble-style model fusion rather than traditional fine-tuning or adapter strategies. Lastly, the dataset Starling-LM-7B-alpha, known for alignment-focused training, contributes to a diverse range of downstream models, including 165 adapters and 18 quantized versions, illustrating its utility in fine-tuning and compression workflows.
One model versus multiple datasets refers to the case when an LLM model is trained with multiple datasets. We observed that DeBERTa-ST-AllLayers-v3.1, a fine-tuned variant of the DeBERTa architecture, takes the top position, having been trained on 116 different datasets. Its adapter-based counterpart, DeBERTa-ST-AllLayers-v3.1bis, also leverages the same number of datasets via adapter-based training, emphasizing modular reuse across tasks.
@@
Dataset Total Fine-tune Adapter Quantization Merges
Mistral-7B-v0.1 1093 300 300 193 300
@@
In addition, models like static-similarity-mrl-mul-v1 and static-similarity-mrl-multilingual are both fine-tuned on 108 datasets, indicating their role in multilingual and multi-task similarity-based retrieval applications. The ModernBERT-base-embed and Llama-3.2-3B-Instruct families show strong dataset diversity as well, with 88 and 87 datasets, respectively, across fine-tuning and quantized variants (e.g., GGUF format). Interestingly, most of these models are fine-tuned, specifically, 8 out of the top 10, highlighting a trend where base models are transformed into task-specific variants through diverse training datasets. This pattern suggests that fine-tuning remains a dominant strategy for adapting base models to downstream tasks across heterogeneous data sources.
Finding #5: Models and datasets exhibit strong bidirectional interdependence, with datasets like Mistral-7B-v0.1 spawning hundreds of models, while models such as DeBERTa-ST-AllLayers-v3.1 leverage diverse datasets to enhance adaptability, highlighting the critical role of dataset-model interactions in advancing AI.
4.6. RQ #6: Dynamic Update Evaluation
@@
Figure 5. The number of changed models and datasets on Hugging Face from June 25 to July 15, 2025.

### v3

@@ 4. Experiment and Finding
  • RQ #4: What are the relationships between models and datasets?
  • RQ #5: What insights can be gained from the dynamic updates?
  • RQ #6: How can HuggingGraph be applied to other platforms?
@@ 4.1. RQ #1: Supply Chain Graph Properties
Graph basics. The collected supply chain graph is a medium-scale directed heterogeneous graph with 402,654 nodes and 462,524 edges as of June 30th, 2025. In particular, there are six different types of nodes, including 28,384 base models, 115,211 fine-tuned, 79,254 adapters, 98,143 quantization models, 13,028 merges, and 68,634 datasets. The average degree is about 1.15, denoting that it is a very sparse graph. Furthermore, we identified substantial metadata missing. As of June 30, 2025, among 1.8 million models, only 50,156 (2.79%) provides a model tree, while $\sim$550K models lack any metadata beyond their names, and nearly another 400K models are empty. Similarly, of the 450K datasets, only 68,634 (15.26%) provide dataset cards, leaving $\sim$380K datasets without any metadata.
@@
Degree distribution shows how node degrees (the number of edges connected to a node) are distributed across the graph. Figure 3 illustrates the indegree and outdegree distribution of the graph. We show not only the total distribution but also the distribution of six types of nodes, including base models, fine-tuned models, adapter models, quantization models, merged models, and datasets.
We observe that the degree distribution in our supply chain graph is heavy-tailed. In particular, the indegree distribution shows a large spread across different categories. The outdegree distribution follows a similar pattern but may differ in specific cases (e.g., adapters seem to have a more restricted degree distribution). This heavy-tailed behavior suggests that most nodes have low degrees, while a few central nodes (hubs) dominate the graph. In particular, the dataset macrocosm-os/images has the highest indegree 550, and dataset Mistral-v0.1 from “mistral AI” has the highest outdegree 1,093. Specifically, the base models act as high-degree hub nodes as they are heavily used by other task-specific models.
@@ 4.2. RQ #2: Supply Chain Structural Analysis
Figure 4. Cumulative distribution function (CDF) of WCC.
@@
Connectivity analysis. We computed weakly connected components (WCCs), which identify a maximal subset of nodes that remain connected when edge directions are ignored (Ji and Huang, 2020; Ji et al., 2020). The total number of WCCs in our supply chain graph is 44,908.
@@
Figure 4 shows the cumulative distribution function (CDF) of the WCC distribution. We observe (i) the largest WCC covers 247,244 nodes, accounting for 61.4% of all the nodes. It reflects the dense interconnections that pervade the ecosystem. This vital element is essential for effective information sharing, resource allocation, and structural support, and is the base of the ecosystem. In the largest WCC, major models are included, such as Gemma-2B, DistilBERT, and GPT-2. (ii) In contrast, the remaining WCCs collectively hold 38.6% of the nodes, with most having 1, 2, or 3 nodes. This indicates a fragmented outer edge characterized by specialized models, rare datasets, or active experimental projects. The prevalence of these small, isolated pieces suggest niche attempts that lack integration with the overall system. For example, zhongqy/RMCBench (benchmarking dataset) and yigit69/bert-base-uncased-finetuned-rte-run_3 (recognizing textual entailment task) remain disconnected due to limited reuse or insufficient metadata.
@@
We also computed strongly connected component (SCC) (Ji et al., 2018), a maximal subgraph in which every node is reachable from every other via directed edges, identifying 398,198 SCCs. Remarkably, only 591 of these are non‑trivial (size ¿ 1), collectively encompassing 5,047 nodes (1.25% of the graph). The largest SCC comprises 478 dataset nodes, among them tree‑of‑knowledge and OpenHermes‑2.5, forming a tightly‑knit cluster. By contrast, the remaining 99.46% of nodes each reside in trivial (size‑1) SCCs.
Community detection identifies node groups with dense internal connections. In LLM supply chains, it reveals aligned subgraphs reflecting reuse patterns and task-specific assets. We apply the Louvain method (De Meo et al., 2011), a greedy algorithm that maximizes modularity to find densely connected communities. Higher modularity indicates stronger intra-community connectivity.
@@
Table 2 summarizes the top-10 communities. (i) We find that each of the top-10 communities achieve a high modularity score of 0.96, indicating strong intra-community connectivity. Collectively, these communities span a wide range of functional domains, underscoring the presence of well-defined, task-aligned clusters within the ecosystem. (ii) The largest community consists of 9,390 nodes and attains a modularity score of 0.96, indicating an extremely cohesive internal structure. It revolves around base models like OLMoE and CausalLM, and general-purpose datasets like prompt-perfect. This suggests a densely connected cluster facilitating widespread reuse and fine-tuning. (iii) Several other communities reflect clear task-based segmentation. For instance, community 2 (7,388 nodes) focuses on solving mathematical problems, with models like qwen2.5_math and datasets like Marco-o1. Similarly, community 6 (4,554 nodes) centers on instruction-tuning, with model MedLlama-3-8B and dataset dpo-mix.
@@
Finding #2: The LLM supply chain graph features a dominant core (61.4% of nodes), while high modularity (0.96) reveals task-aligned, semantically coherent communities amid a fragmented periphery.
4.3. RQ #3: Supply Chain Analysis of Models
@@
Base model impact. We would like to understand the impact of base models. Here, we quantify the impact of a base model as the number of task-specific models that depend on it. The more dependencies, the larger the impact it has. We start with a base model and perform forward analysis by computing BFS following the outgoing edges. This leads to a forward subgraph, which denotes all the models that depend on the base model, including fine-tuned, adapted, quantized, or merged models.
Table 3 shows the top-10 base models sorted by the forward subgraph size, which is the number of impacted task-specific models. We find (i) a base model can significantly impact the LLM supply chain ecosystem. For example, Llama-3.1-8B is a base model from Meta used for efficient text generation, code assistance, and research (He et al., 2024b). Due to its relatively small size, which allows for deployment in resource-constrained environments, making advanced AI accessible to broader stakeholders (Grattafiori et al., 2024). It has generated up to 7,544 models, including 1,710 fine-tuned versions, 1,542 adapters, 3,473 quantizations, and 1,693 merged models tailored to specific tasks. (ii) For fine-tuning, the base model Mistral-7B-v0.1 has been fine-tuned the most, totaling 2,105. It is a faster, lighter Mistral model trained by Mistral AI with grouped-query and sliding-window attention, enabling efficient text generation, NLP, and code assistance on consumer hardware for low-latency tasks (Thakkar and Manimaran, 2023).
@@
Llama-3.1-8B 7,544 1,710 1,542 3,473 1,693 25
Mistral-7B-v0.1 6,744 2,105 2,187 1,435 1,254 27
Qwen2.5-7B 6,733 1,972 1,764 2,516 1,132 11
Meta-Llama-3-8B 5,633 967 1,511 2,220 1,967 21
Llama-3.1-70B 4,063 698 281 2,075 2,519 11
Qwen2.5-32B 3,909 1,086 158 2,311 1,049 12
Qwen2.5-1.5B 3,645 1,300 1,290 949 248 8
Qwen2.5-0.5B 3,521 1,669 1,006 810 46 11
Qwen2.5-14B 3,362 726 411 1,880 1,166 15
Meta-Llama-3-8B-Instruct 3,118 640 405 1,394 1,305 34
@@
Task-specific model analysis. Given a task-specific model, we want to understand how it evolves, e.g., what other models it relies on. To achieve that, for each model, we perform a backward analysis by running BFS following the incoming edges. To that end, the derived subgraph shows the models it relies on.
Table 4 shows the top-10 task-specific models sorted by the backward subgraph size, which is the number of models they rely on. We make two interesting observations. (i) A fine-tuned model, command-r-1-layer, illustrates the depth and complexity of transformations in the LLM supply chain. This model operates in bfloat16 (BF16) precision for efficient text generation and natural language understanding (Pignatelli et al., 2024), originates from the base model c4ai, and has undergone extensive lineage evolution before reaching its final form. Specifically, it depends on 40 upstream artifacts, including 39 other fine-tuned models, and spans 39 transformation levels in its backward lineage chain, as detailed in Table 4. (ii) We observe that adapters are mainly used for lightweight fine-tuning and merges for model integration, but task-specific models like command-r-1-layer, as optimized standalone derivatives, do not evolve from adapters or merges in their backward lineage (Pignatelli et al., 2024).
4.4. RQ #4: Supply Chain of Models and Datasets
@@
This research question explores the interconnections between models and datasets from dual perspectives, including one dataset versus multiple models and one model versus multiple datasets.
@@
One dataset versus multiple models refers to the case when a single dataset is used to train multiple models. Table 5 shows the top-10 datasets based on the number of models trained on them. In particular, (i) Mistral-v0.1 takes the leading position and is a widely adopted open-source dataset known for its strong performance in general-purpose language understanding and generation tasks. It has been used to train 1,093 models, including 300 fine-tuned variants, 300 adapters, 193 quantized models, and 300 merged models, highlighting its broad adoption across diverse model derivation strategies.
@@
Dataset Total Fine-tune Adapter Quantization Merge
Mistral-v0.1 1,093 300 300 193 300
@@
(ii) The dataset TinyLlama-1.1B-v1.0 has been used to train 728 models, featuring 300 fine-tuned variants and 300 adapters. Similarly, open_llama_3b, an open-access dataset of Llama, supports 285 adapter-based models, indicating a preference for lightweight, modular adaptation. (iii) Further, MAmmoTH2-7B stands out with 210 merged models, showcasing its role in ensemble-style model fusion rather than traditional fine-tuning or adapter strategies.
One model versus multiple datasets refers to the case when an LLM model is trained with multiple datasets. Table LABEL:tab:dataset_model_2 shows the top-10 models ranked by the number of datasets used for training. We observe that DeBERTa-ST-AllLayers-v3.1, a fine-tuned variant of the DeBERTa model, takes the top position, having been trained on 116 different datasets. Its adapter-based counterpart, DeBERTa-ST-AllLayers-v3.1bis, also leverages the same number of datasets via adapter-based training, emphasizing modular reuse across tasks.
Table 6. Top-10 models sorted by # of training datasets.
Model
Model Type
# of training dataset
DeBERTa-ST-AllLayers-v3.1 Fine tune 116
DeBERTa-ST-AllLayers-v3.1bis Adapters 116
static-similarity-mrl-mul-v1 Fine tune 108
static-similarity-mrl-multilingual Fine tune 108
ModernBERT-base-embed Fine tune 88
Llama-3.2-3B-Instruct Fine tune 87
Llama-3.2-3B-Instruct-GGUF Quantization 87
DavidLanz-3.2-3B-Instruct Fine tune 87
static-retrieval-mrl-en-v1 Fine tune 79
XLMRoBERTaM3-CustomPoolin Fine tune 72
@@
Figure 5. The number of changed datasets and models on Hugging Face from June 25 to July 15, 2025.
In addition, models like static-similarity-mrl-mul-v1 and static-similarity-mrl-multilingual are both fine-tuned on 108 datasets, indicating their roles in multilingual and multi-task similarity-based retrieval applications. Interestingly, most of these models are fine-tuned, specifically, 8 out of the top-10, suggesting that fine-tuning remains a dominant strategy for adapting base models to downstream tasks across heterogeneous data sources.
Finding #4: Models and datasets exhibit strong bidirectional interdependence, with datasets like Mistral-v0.1 spawning hundreds of models, while models such as DeBERTa-ST-AllLayers-v3.1 leverage different datasets to enhance adaptability, highlighting the critical roles of dataset-model interactions in advancing AI.
4.5. RQ #5: Dynamic Update Evaluation